A manufacturer of light bulbs produces bulbs that last a mean of 950 hours with a standard deviation of 120 hours. What is the probability that the mean lifetime of a random sample of 10 of these bulbs is less than 900 hours? what is the sampling distribution of the sample means


1
Expert's answer
2022-03-16T17:06:48-0400

"P(\bar X<900)=P(Z<\frac{900-950}{\frac{120}{\sqrt{10}}})=P(Z<-1.32)=0.0934."

Sampling distribution: "\mu_{\bar X}=950,\sigma_{\bar X}=\frac{120}{\sqrt{10}}=37.95."